Equipment was purchased on January 5, Year 1, at a cost of $90,000. The equipment had an estimated useful life of eight years and an estimated residual value of $8,000. After using the equipment for three years, the useful life was revised to a total of 10 years and the residual value was reduced to $2,004 Determine the straight-line depreciation expense for Year 4 and the following years.
Answer #1

Depreciation for year 4 and following years= $8178

Working

Straight line Method for 3 years

A

Cost

$ 90,000.00

B

Residual Value

$    8,000.00

C=A - B

Depreciable base

$ 82,000.00

D

Life [in years ]

                      8

E=C/D

Annual SLM depreciation

$ 10,250.00

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$    90,000.00

$ 10,250.00

$ 79,750.00

$   10,250.00

2

$    79,750.00

$ 10,250.00

$ 69,500.00

$   20,500.00

3

$    69,500.00

$ 10,250.00

$ 59,250.00

$   30,750.00

Depreciation schedule for year 4

Straight line Method

A

Book value at the end of Year 3

$ 59,250.00

B

Residual Value

$    2,004.00

C=A - B

Depreciable base

$ 57,246.00

D

Life [in years left ] (10-3)

                      7

E=C/D

Annual SLM depreciation

$    8,178.00

Year

Book Value

Depreciation expense

4

$    59,250.00

$    8,178.00